Tim and Judy mix two kinds of feed for pedigreed dogs. They wish to make 27 pounds of feed worth $0.26 per pound by mixing one kind worth $0.17 per pound with another worth $0.36 per pound. How many pounds of the cheaper kind should they use in the mix?
Answer by mananth(16949)

one kind worth $0.17 per pound let this be x lbs.
with another worth $0.36 per pound. this will be 27-x lbs
27 pounds $0.26 per pound
..
0.17*x+0.36(27-x)=27*0.26
0.17x+9.72-0.36x= 7.02
-0.19x = -9.72+7.02
-0.19x=-2.70
x= -2.70/-0.19
= 14.2 lbs.